Lance Whittingham IV specializes in buying deep discount bonds. These represent bonds that are trading at well below par value. He has his eye on a bond issued by the Leisure Time Corporation. The $1,000 par value bond pays 7 percent annual interest and has 16 years remaining to maturity. The current yield to maturity on similar bonds is 10 percent. Use Appendix B and Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods. a. What is the current price of the bonds? (Do not round intermediate calculations. Round your final answer to 2 decimal places. Assume interest payments are annual.)

a. What is the current price of the bonds? (Do not round intermediate calculations. Round your final answer to 2 decimal places. Assume interest payments are annual.)

b. by what percent will the price of the bonds increase between now and maturity?
Answer #1

GIVEN IN THE PROBLEMS

BOND PAR VALUE= $1,000

YTM(YIELD TO MATURITY) = 10%

MATURITY PERIOD = 16 YEARS

ANNUAL INTEREST RATE = 7%

PMT/INTEREST PER PERIOD = 1000 * 7% = 70 $

a) TO FIND THE CURRENT PRICE OF BONDS

PRESENT VALUE FUNCTION IS UTILIZED = PV(RATE,NPER,PMT,FV,TYPE)

= PV(10%,16,70,1000,0) * -1

CURRENT PRICE OF BOND = $765.29

B) PERCENTAGE INCREASE

= (1000/765.29)-1 = .3067 * 100 = 30.67%

NOTE: THE COMPOUNDING PERIOD HERE IS ONE(1)